TCCON > Sites > Wollongong > Instrument History > Wollongong TCCON Event Log

Wollongong TCCON Event Log

Table of contents
No headers

 

 
EventDate[yyyy/mm/dd] EventNumber Event
2008/10/08 0 Start NIR-DC acquisition
2009/01/07 1 Change tracker mirrors
2009/02/20 2 Tracker 45 input mirror replaced
2009/03/16 3 Beamsplitter clamp loose (discovered later)
2009/04/08 4 Realign laser and IR
2009/04/08 5 Align tracker
2009/04/17 6 Tracker input mirror replacement
2009/04/29 7 Tracker main mirror replacement
2009/05/18 8 Tracker settled
2009/05/25 9 IR realignment
2009/07/25 10 Fixed CC alignment
2009/08/20 11 Phase resolution reset
2009/08/24 12 New perspex source cover
2009/08/31 13 Reduce intensity with gauze
2009/09/24 14 Recentre tracker pickoff mirror
2009/11/10 15 Clean tracker mirrors
2009/11/12 16 Minor laser alignment
2009/11/04 17 Set CaF2 beamsplitter gain High
2009/12/02 18 Pressure sensor cal changed
2009/12/10 19 Revert pressure cal
2010/01/19 20 Change input window to KCl  realign
2010/02/09 21 Remove input window gauze
2010/02/26 22 Change input window to repolished KBr
2010/03/26 23 Tweak tracker
2010/03/31 24 Pressure calibration added to sunrun
2010/04/15 25 Overmodulation  reason unclear
2010/04/20 26 Replace tracker 45deg input mirror
2010/06/13 27 Install input window gauze
2010/06/15 28 Gauze removed
2010/06/30 29 Corner cube shear alignment
2011/03/04 30 Tweak tracker input
2011/07/22 31 Laser board ECL-04. Adjust ghosts.
2012/02/24 32 Tracker realignment
2012/10/02 33 Full realignment by fringes
2012/10/29 34 Disk died  reinstall OS
2012/12/07 35 Input beam selector failure
2012/12/18 36 Beam selector fixed
2013/02/19 37 Clean tracker mirrors
2013/05/15 38 Ghost minimisation  minor APT adjustment
2013/09/30 39 KBr beamsplitter gain low
2013/10/21 40 Change tracker thresholds
2013/10/11 41 output beam selector failed
2013/11/05 42 output beam selector fixed
2014/03/31 43 Clean tracker mirrors
2014/08/03 44 Laser dead
2014/08/19 45 Laser replaced (REO)  Win7 installed
2014/10/23 46 Tracker mirrors cleaned
2015/02/17 47 Tracker mirrors cleaned
2015/03/18 48 5V power supply low
2015/04/30 49 Laser replaced (SIOS)
2015/05/19 50 Evacuate detectors
2015/10/19 51 Minor realignement and ghost measurements
2015/12/18 52 Rogue cable may be in the solar beam
2016/01/12 53 Vacuum Failure  vented measurements
2016/01/14  54 Replaced PS16 (M15E-PS(I0956)) board
2016/01/17 55 Resumed evacuated measurements (small pump)
2016/04/06 56 Beam selector motor failure  manual beam selection.
2016/04/21  57 Cleaned tracker mirrors
2016/05/10 58 Measurements without HCl cell
2016/05/23 59 Motor drive replaced  HCl cell returned
2016/05/25  60 Red rogue cable in the solar beam
2016/07/01  61 Rogue cable within ca 40% of solar beam
2016/08/22 62 Adixen ACP 40 Vac pump reinstalled
2016/11/11 63 Rogue cable in solar beam path. S-G shifts increased x3.
2016/11/21 64 Rogue cable in solar beam path removed.
2017/01/09  65 Black rogue cable in the solar beam
2017/02/06 66 Rogue tracker cable in the solar beam
2017/04/01 67 Start of solar tracker jamming
2017/05/01 68 Cleaned tracker mirrors. Some flaking on 45 deg entrance miror
2017/07/04 69 Possum chewed tracker wires
2017/08/04 70 Higher S-G shifts due to humidity sensor cable loosened by possum

 

Tag page